The vapour pressure of pure benzene and toulene are 160 and 60 torr respectively. The mole fraction of toulene bis vapour phase in contact with equimolar solution of benzene and toulene is -
A)0.50
B)0.6
C)0.27
D)0.73

Grade:12

1 Answers

Arun
5 years ago
Pb = 0.5 * 160 = 80 Torr
 
Now Ptotal = 0.5 * 160 + 0.5 * 60 = 110 Torr
 
hence
yb = Pb/ Ptotal = 80/110 = 0.727
 
Hope it helps